OpenStudy (gabbyalicorn):

A music shop has 5 different brands of guitar strings. The 5 different brands sell for $16, $20, $15, $11, and $13. What is the mean absolute deviation of the prices? Round to the nearest cent, if necessary. @data_LG2

OpenStudy (gabbyalicorn):

The mean is 19. Now I need help with the rest.

OpenStudy (gabbyalicorn):

@LegendarySadist

OpenStudy (anonymous):

Do you know the formula for MAD? Here it is how it looks like:  It will be a long process... Steps: 1. You already find the mean :) 2. Get the difference of the mean from each value 3. Get the sum of each differences 4. Then divide it by the number of the data set.

OpenStudy (gabbyalicorn):

so difference = subtracting right? but what do i subtract it (19) with?

OpenStudy (anonymous):

wait, for the mean are you sure you'll get 19?

OpenStudy (gabbyalicorn):

yeppers. :)

OpenStudy (anonymous):

what is 16+20+15+11+13 ?

OpenStudy (gabbyalicorn):

oh, i got 15

OpenStudy (anonymous):

yep :) Now we got the same thing. Step number two: \(\sf x-x_{mean}\) For example, you have x= 16 and xmean is 15, so 16-15 is 1. Do this for x=20, x= 15, x=11 and x=13 and tell me what you get :)

OpenStudy (gabbyalicorn):

20 - 15= 5 11- 13= 2

OpenStudy (anonymous):

hmm.. given data-mean so it will be like 16-15 = 1 20-15 = ? 15-15= ? 11-15 =? 13-15 = ?

OpenStudy (gabbyalicorn):

oh, i see sorry

OpenStudy (anonymous):

no worries :)

OpenStudy (gabbyalicorn):

20-15 = 5 15-15= 0 11-15 = 4 13-15 = 2

OpenStudy (anonymous):

right, since we are getting the absolute values, we'll ignore the negative differences:) So far we have the differences: 1, 5, 0, 4, 2 now let's move to step 3: Add the values you get from step 2 :)